OpenStudy (anonymous):

How do you graph exponential functions? I'm really confused on how to solve them. Can you please help? For instance y = 4x. It looks so simple. But I'm so confused.

jimthompson5910 (jim_thompson5910):

one way to graph any function is to plot a series of points

jimthompson5910 (jim_thompson5910):

to get any point, you plug in an x value to find a corresponding y value

jimthompson5910 (jim_thompson5910):

so for instance, say x = 1 if x = 1, then y = 4^x y = 4 which means the point (1,4) is on the curve

jimthompson5910 (jim_thompson5910):

do this with enough x values and you get enough points with enough points, you can draw a curve through them all (the more points you have, the more accurate the graph is)

OpenStudy (anonymous):

So, I could plug in 1, 2, 3, 4, and so forth until I have a line and I can plot it?

jimthompson5910 (jim_thompson5910):

it won't be a straight line though

jimthompson5910 (jim_thompson5910):

\(\large y = 4^x\) is NOT a straight line (it's a curve that bends upwards)

OpenStudy (phi):

two other factoids you need to remember: \[ 4^0 = 1 \] and \[ 4^{-m} = \frac{1}{4^m} \] you need this one for negative x. for example, when x=-2 \[ 4^{-2} = \frac{1}{4^2} = \frac{1}{4\cdot 4}= \frac{1}{16}\]